When a force of 0.1 N is applied, the spring is stretched by 1.5 cm. The spring is cut into three parts and one part is stretched by 3 cm. Find the force required for doing so.

 

Grade:12th pass

2 Answers

vishal akula
5 years ago
when spring is cut t, k of spring changes
so use realtion length of springis directly proportional to lenth and find k 
use F=kx relation and find f
vishal akula
5 years ago
given intially F=0.1N for to strech by 1.5cm that is k=F/x=100/15
for l lenth k=100/15 that is for l/3,, k= 300/15=20Nm
that one part k=20Nm to strech it by 3cm=0.03m F=0.6N